OpenStudy (anonymous):

A mass m = 4.6 kg hangs on the end of a massless rope L = 2.16 m long. The pendulum is held horizontal and released from rest. 1)How fast is the mass moving at the bottom of its path?2)What is the magnitude of the tension in the string at the bottom of the path?

OpenStudy (anonymous):

Find the velocity my using the conservation of energy mgh=1/2mv^2. At the bottom, the mass is turning in a circle with tension pointing up towards the center of the circle and weight pointing down. Use the centripetal force requirement mv^2/r=T-W and solve for T. Hope this helps.

OpenStudy (anonymous):

If the maximum tension the string can take without breaking is Tmax = 384 N, what is the maximum mass that can be used? (Assuming that the mass is still released from the horizontal and swings down to its lowest point.

OpenStudy (anonymous):

Does this require to solve for m using the quadratic formula? Little lost on how to reach m.

OpenStudy (anonymous):

Use the quadratic formula? where is mass ever squared in this question?